if a product normally retails for 40.00 and we give you a $6 discount , what percentage discount are we givig you ?
Answer by missgigid(8)

$6/$40=X/100
$40X=600
$40X/40=600/40
X=15%
15% IS THE PERCENTAGE OF THE DISCOUNT BEING GIVEN.